# A Deep Learning Framework for Disentangling Triangle Singularity and Pole-Based Enhancements

### v2

@@ Abstract
Enhancements in the invariant mass distribution or scattering cross-section are usually associated with resonances. However, the nature of exotic signals found near hadron-hadron thresholds remain a puzzle today. In fact, a purely kinematic triangle mechanism is also capable of producing similar structures, but do not correspond to any unstable quantum state. In this paper, we report for the first time, that a deep neural network can be trained to distinguish triangle singularity from pole-based enhancements. We also identify the type of dynamic pole structure that can be misidentified as triangle enhancement. We apply our method to confirm that the $P_{\psi}^{N}(4312)^{+}$ state is not due to a single triangle singularity, and is more favored towards a pole-based interpretation based solely on pure line-shape analysis. Lastly, we argue how our method can be used as a model-selection framework to help in classifying other exotic hadron candidates.
@@ I Introduction
In addition to compact and hadronic molecules, some signals do not correspond to unstable quantum states. These signals are generally referred to as cusps and are typically associated with the kinematics of reactions Guo et al. (2020). Due to the branch-point singularity associated with the opening of a new channel, one expects a sudden change in the line shape of an $s$-wave amplitude at the threshold. This change in the line shape is further enhanced by the presence of poles located even in the distant Riemann sheet. The problem with cusps is that they can be misinterpreted as quantum states due to the uncertainty present in the experimental data. It therefore warrants more careful analysis to discriminate them from true resonances Guo et al. (2015a). Distinguishing different origins of signals, especially near thresholds, is one of the long-standing issues in hadron spectroscopy.
@@
One of the novel approaches in analyzing line shapes is through the use of machine learning techniques Albaladejo et al. (2022); He et al. (2023); Boehnlein et al. (2022). In particular, the problem of line shape interpretation can be phrased as a classification problem where the input corresponds to line shape and the output is the interpretation or the origin of enhancements. Once the classification problem is unambiguously defined, a deep neural network (DNN) can be trained to map the input line shape space to the output interpretation space. This mapping is made possible because a trained DNN can be considered as a universal approximator Hornik et al. (1989); Schmidhuber (2015). Such line shape analysis framework was first utilized in Sombillo et al. (2020) where a neural network was developed to distinguish between a bound and a virtual state enhancement at the nucleon-nucleon threshold. The method was later applied to the pion-nucleon system in Sombillo et al. (2021) where a DNN is trained to identify the distribution of poles in different Riemann sheets. The expected next step is to use the machine learning technique to analyze exotic signals. This attempt was first done in Ng et al. (2022) where a DNN was trained using a coupled-channel effective range approximation to identify the origin of the $P_{\psi}^{N}(4312)^{+}$ signal. It was further demonstrated in the same study that the machine learning technique conforms with the result of the conventional analysis made in Fernández-Ramírez et al. (2019). Further investigations of other exotic signals using machine learning are done in Liu et al. (2022); Zhang et al. (2023); Chen et al. (2023).
@@
Existing machine learning studies have been focused entirely on the conventional interpretations of exotic signals and have not yet considered the possible role of kinematical effects. One cannot simply rule out the kinematic interpretation when it has been shown to also produce resonance-like enhancements near two-hadron thresholds Guo et al. (2020, 2015a). A more general approach would be to train a DNN to recognize and distinguish line shapes arising from the dynamic pole structure and the kinematical effects in a reaction Liu et al. (2016a). This way, it can be used to identify the origin of signals without any particular biases. One such kinematical mechanism prominently used in the study of exotic signals is the triangle singularity.
@@
First proposed by Landau in 1960 Landau (1959), the triangle singularity is a kinematic re-scattering process involving three intermediate particles, such that the decay amplitude has a singularity close to the physical region which produces an enhancement in the invariant mass distribution of the final state Guo et al. (2020). This process is represented by a triangle diagram, illustrated in Figure 1, involving an initial state $A$ decaying into final states $B$ and $C$, with intermediate particles 1, 2, and 3. During the process, particle $A$ decays into particles 1 and 2, then particle 1 decays into particles 3 and $B$. Particle 3 then catches up to particle 2 along the same direction and collides to form particle $C$ in the final state. For it to produce enhancements in the physical region, it is required that the interaction at all vertices be classical processes and all the intermediate particles be real propagators (on their mass shell with real momenta) Coleman and Norton (1965).
@@
The triangle singularity regained prominence in the community when it was recently used to explain some newly observed exotic signals Wu et al. (2012, 2013); Mikhasenko et al. (2015); Aceti et al. (2016); Szczepaniak (2015); Nakamura (2020). In particular, it was suggested that the exotic hidden-charm pentaquark state $P_{c}(4450)^{+}$ observed by the LHCb collaboration in 2015 Aaij et al. (2015) could be the manifestation of a $\Lambda(1890)\chi_{c1}p$ triangle loop in the $\Lambda_{b}^{0}\to J/\psi\,pK^{-}$ decay Guo et al. (2015b); Liu et al. (2016b); Bayar et al. (2016); Mikhasenko (2015). The renewed interest in studying this mechanism is due to the similarities between the enhancements produced by a triangle singularity and that of a true hadronic resonance, which once again stresses the issues of ambiguity in interpreting experimental data. Since then, the nature of triangle singularity and its application in analyzing exotic signals has been extensively studied. For a complete review, see Guo et al. (2020).
Most recently, an updated analysis of the decay was reported by LHCb in 2019, in which the narrow pentaquark state $P_{\psi}^{N}(4312)^{+}$ (then called the $P_{c}(4312)^{+}$), together with the two-peak structure of the $P_{c}(4450)$ resonance, was discovered Aaij et al. (2019). In their analysis of these signals, the triangle mechanism was ruled out as a reasonable interpretation for the $P_{\psi}^{N}(4312)^{+}$ and $P_{c}(4440)^{+}$ states due to lack of proper hadron-rescattering thresholds in the internal triangle loop diagrams. It, however, remains a plausible hypothesis for the $P_{c}(4457)^{+}$ state, requiring more thorough amplitude analysis to establish a definitive interpretation. Further studies have also pointed out the possibility of a more general double triangle mechanism producing similar structures Nakamura (2021).
In this study, we apply machine learning techniques to analyze the various interpretations of the $P_{\psi}^{N}(4312)^{+}$ state, including the relevant kinematical effects. We develop, for the first time, a deep neural network capable of distinguishing triangle singularity from pole-based enhancements. Our goal is to apply the trained DNN on the $P_{\psi}^{N}(4312)^{+}$ state and validate if the triangle singularity interpretation can indeed be confidently ruled out. In doing so, we demonstrate the effectiveness of our novel model-selection framework and its significance as a possible tool for classifying exotic hadron candidates.
@@ II Triangle Singularity
where $P$, $k$, $q$ are the momenta of the particles as labelled in Figure 1. The masses $m_{1}$, $m_{2}$, and $m_{3}$ correspond to the masses of the intermediate particles, and $\epsilon$ pertains to the their decay width in the internal diagram.
Since the intermediate particles must be real propagators, one can apply approximations as discussed in Guo et al. (2020), and rewrite the integral as
@@
where $q=|\vec{q}|$ is now the magnitude of the 3-momentum vector and $P^{0}$ is the total energy of the initial state. This simplified form is sufficient in the investigation of the underlying singularities that will emerge from the triangle loop. Since particles $m_{1}$ and $m_{2}$ are on-shell, the integrand has a singularity at values of $q$ satisfying $P^{0}=\sqrt{m_{1}^{2}+q^{2}}+\sqrt{m_{2}^{2}+q^{2}}-i\epsilon$, which corresponds to the left blue cut appearing in Figure 1. The function $f(q)$ in equation (2) is the angular part of the volume integral, which is given by
$f(q)=\displaystyle\int_{-1}^{1}\dfrac{dz}{E_{C}-\omega(q)-\sqrt{m_{3}^{2}+q^{2}+k^{2}+2qkz}+i\epsilon},$
where $E_{C}=(m_{A}^{2}-m_{B}^{2}+m_{C}^{2})/2m_{A}$ is the energy of particle $C$ in the rest frame of $A$, $\omega(q)=\sqrt{m_{2}^{2}+q^{2}}$, and $k=\sqrt{\lambda(m_{A},m_{B},m_{23})}/2m_{A}$, with $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2yz-2xz$. $m_{23}$ is the invariant mass of the final state. By straightforward substitution, one can calculate the integral in equation (3) and find that it contains a logarithmic expression given by
@@
where $u_{\pm}=E_{C}-\omega(q)-\sqrt{m_{3}^{2}+q^{2}+k^{2}\pm 2qk}$. The logarithmic singularity at $q$ satisfies $u_{\pm}=0$, corresponding to the right red cut in Figure 1.
To ensure the convergence of the integral, we may include a form factor with an expression given by $\Lambda^{2}/(\Lambda^{2}+q^{2})$, where we define $\Lambda$ as some arbitrary cut-off parameter which can be related to the range of hadronic interaction Nakamura (2020).
Essentially, we find that with equations (2) and (4), the integral $I$ can now be expressed as a simple function of all the involved particle’s masses. In other words, the underlying singularities are completely dependent on kinematical variables alone and shall not require further analysis of the interactions within the system. It is, however, sensitive to these variables, meaning that the enhancement would not occur if certain kinematic conditions are not satisfied. Nevertheless, we have arrived at a useful result that allows us to directly compute for $|I|^{2}$ and construct the line shape from any triangle diagram once we determine the involved masses.
@@
which consequently establishes a constraint for the masses allowed in the triangle loop. These prescribed ranges shall become relevant when choosing the appropriate triangle loops capable of producing resonance-like peaks in a given reaction. They also imply that by adjusting the masses of the intermediate particles, the position of the resulting peak in the invariant mass may be controlled. This shall be our main guiding principle for generating multiple TS line shapes as training data for the DNN, which shall be discussed in section IV.
The $P_{\psi}^{N}(4312)^{+}$ structure, with a measured signal at $4311.9\pm 0.7$ MeV, can be observed just below the $\Lambda_{c}^{+}\bar{D}^{*0}$ threshold. From the analysis performed by the LHCb Aaij et al. (2019), a triangle loop of $\Lambda_{c}^{+}D_{s}^{**-}\bar{D}^{*0}$ can be constructed where the mass of some hypothetical $D_{s}^{**-}$ meson is set to 3288 MeV, such that the resulting line shape can fit the experimental data with a peak at 4312 MeV. The triangle diagram is shown in Figure 3(a).
Using the Particle Data Group (PDG) values Workman et al. (2022) of the hadron masses involved in the triangle diagram, one can compute the integral in equation (2) and plot the $|I|^{2}$ contribution as a function of the invariant mass of the final state, which in this case is $J/\psi p$. Note that for demonstration purposes, we use convenient, albeit unrealistic, values of $\epsilon=1$ MeV and $\Lambda=3000$ MeV to evaluate the integral. Furthermore, we can approximate a picture of a resonance using the Breit-Wigner parametrization given by
@@
where $E$ is the scattering energy, $M$ and $\Gamma$ are the mass and decay width of the unstable state. For both of these amplitudes, we multiply a projected phase space of the $J/\psi p$ invariant mass and some arbitrary high-order degree polynomial, following the expression
@@
where $\rho(\sqrt{s})$ is the phase space factor, $F(\sqrt{s})$ is the constructed amplitude (either Breit-Wigner or the triangle integral $I$), and $B(\sqrt{s})$ is the additional polynomial background Fernández-Ramírez et al. (2019). The $dN/d\sqrt{s}$, corresponding to the resulting intensity, achieves a better fit of the experimental data by optimizing the parameters in $F(\sqrt{s})$ and $B(\sqrt{s})$. The resulting picture is shown in Figure 2 where a resonance-like peak at around 4312 MeV is achieved by both line shapes. Since they both fit within the uncertainties (error bars) of the experimental data, further analysis is required to provide an unambiguous discrimination between the two vastly different mechanisms and interpretations yet similar line shapes.
@@
Although this triangle picture fits the experimental data well, it was argued to be an unlikely origin of the enhancement due to some limitations of the mechanism involved Aaij et al. (2019). First, it lacks an appropriate hadron-rescattering threshold, since the $D_{s}^{**-}$ hadron in the loop is only some hypothetical charmed meson that does not exist. Second, for a good fit to be achieved, the mechanism requires unrealistic parameters to be used. When the decay width $\epsilon$ is set to some higher and more plausible value, say 50 MeV, the enhancement from the triangle singularity is no longer prominent and falls weakly into the background. This is why a pole-based interpretation is favored for the $P_{\psi}^{N}(4312)^{+}$ state.
@@
Figure 3: Different mechanisms for the $\Lambda_{b}^{0}\to J/\psi\,pK^{-}$ reaction that produce a peak at 4312 MeV: a) triangle loop $\Lambda_{c}^{+}D_{s}^{**-}\bar{D}^{*0}$, b) triangle loop $J/\psi\Lambda^{*}p$, c) $P_{\psi}^{N}(4312)^{+}$ resonance, and d) coupled-channel decay.
@@
Furthermore, there are other triangle diagrams that can produce peaks in the $J/\psi p$ invariant mass Liu et al. (2016b); Bayar et al. (2016). For instance, a $D_{s}^{**-}\Lambda_{c}^{+}\bar{D}^{0}$ triangle loop can generate enhancements close to the $P_{\psi}^{N}(4312)^{+}$ state as well. In Bayar et al. (2016), they considered multiple diagrams where a general $\Lambda^{*}$ hyperon is assigned as particle 1 (which can couple to $K^{-}p$), a $c\bar{c}$ meson for particle 2, and $p$ for particle 3. For each $c\bar{c}$ candidate, they used equations (5) and (6) to obtain the relevant range of the hyperon mass and the range of the resulting peak. Further analysis of the branching ratios led them to argue the plausibility of each diagram. Among these, if we select the $\Lambda^{*}J/\psi p$ loop with the mass of $\Lambda^{*}$ within [2151, 2523] MeV, a peak can be achieved within the energy region of [4035, 4366] MeV. The triangle diagram is illustrated in Figure 3(b). This is relevant to the $P_{\psi}^{N}(4312)^{+}$ since the signal lies within the allowed energy region of the mechanism. However, upon further inspection, the last two intermediate particles $J/\psi p$ are similar to the two final states in the mechanism. This sort of process must follow the Schmid theorem, which states that if particles 2 and 3 in the internal triangle diagram rescatters to similar final states 2+3, then the TS will simply be absorbed by the $s$-wave of a tree-level diagram and will not produce observable enhancements Schmid (1967). In other words, this mechanism may also be ruled out if any manifestation of the triangle diagram is argued to not be detectable. However, it has been shown in Anisovich and Anisovich (1995) that if one considers an inelastic scattering process in the final vertex, the Schmid theorem becomes irrelevant as the triangle may still be present under coupled-channel effects. Moreover, if a realistic non-zero decay width $\epsilon$ is used in the triangle, the strength of its signal is found to be comparable to that of the tree-level diagram and will also no longer be negligible Debastiani et al. (2019). In any case, we consider both the triangle diagrams in Figures 3(a) and 3(b) in our study as we simply aim to simulate prominent line shapes from triangle mechanisms to investigate their contributions regardless if they are unrealistic or unphysical. These line shapes will be used to train the DNN. One may argue that doing this allows us to possibly “trick” our machine and really test the limits of its abilities. It is up to the DNN to distinguish realistic from unrealistic mechanisms based solely on a pure line-shape analysis. Our end goal is to assess if the DNN may rule out these unrealistic triangle diagrams in favor of the dynamic pole structure for the $P_{\psi}^{N}(4312)^{+}$ state. Therefore, it follows that we take into account both these two triangle mechanisms in our study. From this point on, the first triangle diagram discussed will be referred to as triangle loop A, and the second as triangle loop B.

### v3

@@ Abstract
Enhancements in the invariant mass distribution or scattering cross-section are usually associated with resonances. However, the nature of exotic signals found near hadron-hadron thresholds remain a puzzle today due to the presence of experimental uncertainties. In fact, a purely kinematical triangle diagram is also capable of producing similar structures, but do not correspond to any unstable quantum state. In this paper, we report for the first time, that a deep neural network can be trained to distinguish triangle singularity from pole-based enhancements with a reasonably high accuracy of discrimination between the two seemingly identical line shapes. We also identify the type of triangle enhancement that can be misidentified as a dynamic pole structure. We apply our method to confirm that the $P_{\psi}^{N}(4312)^{+}$ state is not due to a triangle singularity, but is more consistent with a pole-based interpretation, as determined solely through pure line-shape analysis. Lastly, we explain how our method can be used as a model-selection framework useful in studying other exotic hadron candidates.
@@ I Introduction
In addition to compact and hadronic molecules, some signals do not correspond to unstable quantum states. These signals are generally referred to as cusps and are typically associated with the kinematics of reactions Guo et al. (2020). Due to the branch-point singularity associated with the opening of a new channel, one expects a sudden change in the line shape of an $s$-wave amplitude at the threshold. This change in the line shape is further enhanced by the presence of poles located even in the distant Riemann sheet. The problem with cusps is that they can be misinterpreted as quantum states due to the uncertainty present in experimental data. It therefore warrants more careful analysis to discriminate them from true resonances Guo et al. (2015a). Distinguishing different origins of signals, especially near thresholds, is one of the long-standing issues in hadron spectroscopy.
@@
One of the novel approaches in analyzing line shapes is through the use of machine learning techniques Albaladejo et al. (2022); He et al. (2023); Boehnlein et al. (2022). In particular, the problem of line shape interpretation can be phrased as a classification problem where the input corresponds to line shape and the output is the interpretation or the origin of enhancements. Once the classification problem is unambiguously defined, a deep neural network (DNN) can be trained to map the input line shape space to the output interpretation space. This mapping is made possible because a trained DNN can be considered as a universal approximator Hornik et al. (1989); Schmidhuber (2015). Such line shape analysis framework was first utilized in Sombillo et al. (2020) where a neural network was developed to distinguish between a bound and a virtual state enhancement at the nucleon-nucleon threshold. The method was later applied to the pion-nucleon system in Sombillo et al. (2021a) where a DNN is trained to identify the distribution of poles in different Riemann sheets. The expected next step is to use the machine learning technique to analyze exotic signals. This attempt was first done in Ng et al. (2022) where a DNN was trained using a coupled-channel effective range approximation to identify the origin of the $P_{\psi}^{N}(4312)^{+}$ signal. It was further demonstrated in the same study that the machine learning technique conforms with the result of the conventional analysis made in Fernández-Ramírez et al. (2019). Further investigations of other exotic signals using machine learning are done in Liu et al. (2022); Zhang et al. (2023); Chen et al. (2023).
@@
Existing machine learning studies have focused entirely on the conventional interpretations of exotic signals and have not yet considered the possible role of kinematical effects. One cannot simply rule out the kinematic interpretation when it has been shown to also produce resonance-like enhancements near two-hadron thresholds Guo et al. (2020, 2015a). A more general approach would be to train a DNN to recognize and distinguish line shapes arising from both the dynamic pole structure and the kinematical effects in a reaction Liu et al. (2016a). This way, it can be used to identify the origin of signals without any particular biases. One such kinematical mechanism prominently used in the study of exotic signals is the triangle singularity.
@@
First proposed by Landau Landau (1959), the triangle singularity is a kinematic re-scattering process involving three intermediate particles, such that the decay amplitude has a singularity close to the physical region which produces an enhancement in the invariant mass distribution of the final state Guo et al. (2020). This process is represented by a triangle diagram, illustrated in Fig. 1, involving an initial state $A$ decaying into final states $B$ and $C$, with intermediate particles 1, 2, and 3. During the process, particle $A$ decays into particles 1 and 2, then particle 1 decays into particles 3 and $B$. Particle 3 then catches up to particle 2 along the same direction and collides to form particle $C$ in the final state. For it to produce enhancements in the physical region, it is required that the interaction at all vertices be classical processes and all the intermediate particles be real propagators—on their mass shell with real momenta Coleman and Norton (1965).
@@
The triangle singularity regained prominence in the community when it was used to explain some newly observed exotic signals Wu et al. (2012, 2013); Mikhasenko et al. (2015); Aceti et al. (2016); Szczepaniak (2015); Nakamura (2020). In particular, it was suggested that the exotic hidden-charm pentaquark state $P_{c}(4450)^{+}$ observed by the LHCb collaboration in 2015 Aaij et al. (2015) could be the manifestation of a $\Lambda(1890)\chi_{c1}p$ triangle loop in the $\Lambda_{b}^{0}\to J/\psi\,pK^{-}$ decay Guo et al. (2015b); Liu et al. (2016b); Bayar et al. (2016); Mikhasenko (2015). The renewed interest in studying this mechanism is due to the similarities between the enhancements produced by a triangle singularity and that of a true hadronic resonance, which once again emphasizes the issues in interpreting ambiguous experimental line shapes. Since then, the nature of triangle singularity and its application in analyzing exotic signals has been extensively studied. For a complete review, see Guo et al. (2020).
In 2019, an updated analysis of the decay was reported by LHCb wherein the narrow pentaquark state $P_{\psi}^{N}(4312)^{+}$ (then called the $P_{c}(4312)^{+}$), together with the two-peak structure of the $P_{c}(4450)$ resonance, was observed with notable statistical significance Aaij et al. (2019). In their analysis of these signals, the triangle mechanism was ruled out as a reasonable interpretation for the $P_{\psi}^{N}(4312)^{+}$ and $P_{c}(4440)^{+}$ states due to lack of proper hadron-rescattering thresholds in their triangle diagrams. It, however, remains a plausible hypothesis for the $P_{c}(4457)^{+}$ state, requiring further amplitude analysis to establish a more definitive interpretation. Other studies have also pointed out the possibility of a general double triangle mechanism capable of producing similar structures Nakamura (2021).
In this study, we apply machine learning techniques to analyze the various mechanisms relevant to the $P_{\psi}^{N}(4312)^{+}$ state, including the kinematic triangle diagram. We develop, for the first time, a deep neural network capable of distinguishing triangle singularity from pole-based enhancements via pure line-shape analysis. We benchmark our framework by applying the trained DNN on the $P_{\psi}^{N}(4312)^{+}$ state and validate if the triangle singularity interpretation can be consistently ruled out. In doing so, we demonstrate the effectiveness of our novel model-selection framework and its significance as a possible tool for classifying exotic hadron candidates.
@@ II Triangle Singularity
where $P$, $k$, $q$ are the momenta of the particles as labelled in Fig. 1. The masses $m_{1}$, $m_{2}$, and $m_{3}$ correspond to the masses of the intermediate particles, and $\epsilon$ pertains to the decay width of the internal diagram.
Since the intermediate particles must be real propagators, one can apply on-shell treatments as discussed in Guo et al. (2020), and rewrite the integral as
@@
where $q=|\vec{q}|$ is now the magnitude of the 3-momentum vector and $P^{0}$ is the total energy of the initial state. This simplified form is sufficient in the investigation of the underlying singularities that will emerge from the triangle loop. Since particles $m_{1}$ and $m_{2}$ are on-shell, the integrand has a singularity at values of $q$ satisfying $P^{0}=\sqrt{m_{1}^{2}+q^{2}}+\sqrt{m_{2}^{2}+q^{2}}-i\epsilon$, which corresponds to the left blue cut appearing in Fig. 1. The function $f(q)$ in equation (2) is the corresponding angular part of the volume integral given by
$f(q)=\displaystyle\int_{-1}^{1}\dfrac{dz}{E_{C}-\omega(q)-\sqrt{m_{3}^{2}+q^{2}+k^{2}-2qkz}+i\epsilon},$
where $E_{C}=(m_{A}^{2}-m_{B}^{2}+m_{C}^{2})/2m_{A}$ is the energy of particle $C$ in the rest frame of $A$, $\omega(q)=\sqrt{m_{2}^{2}+q^{2}}$, and $k=\sqrt{\lambda(m_{A},m_{B},m_{23})}/2m_{A}$, with $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2yz-2xz$. The $m_{23}$ parameter is the invariant mass of the final state. By straightforward substitution, one can calculate the integral in equation (3) and find that it contains a logarithmic expression given by
@@
where $u_{\pm}=E_{C}-\omega(q)-\sqrt{m_{3}^{2}+q^{2}+k^{2}\mp 2qk}$. The logarithmic singularity at $q$ satisfies $u_{\pm}=0$, corresponding to the right red cut in Fig. 1.
To ensure the convergence of the integral, we may include a form factor with an expression given by $\Lambda^{2}/(\Lambda^{2}+q^{2})$, where we define $\Lambda$ as some arbitrary cut-off parameter related to the range of hadronic interaction Nakamura (2020).
Essentially, we find that with equations (2) and (4), the integral $I$ can now be expressed as a simple function of all the involved particle’s masses. In other words, the underlying singularities are completely dependent on kinematical variables alone and do not involve any dynamical interactions within the system. It is, however, sensitive to these variables, meaning that the enhancement would not occur if certain kinematic conditions are not satisfied. Still, we have arrived at a useful result that allows us to directly compute for intensity $|I|^{2}$ and construct the line shape for any triangle diagram once we determine the involved masses.
@@
which consequently establishes a constraint for the masses allowed in the triangle loop. These prescribed ranges shall become relevant when choosing the appropriate triangle loops capable of producing resonance-like peaks in a given reaction. Another implication is that one may easily control the position of the resulting peak in the invariant mass by simply adjusting the masses of the intermediate particles. We utilize this as our guiding principle for generating multiple TS line shapes as training data for the DNN. This is discussed in section IV.
The $P_{\psi}^{N}(4312)^{+}$ structure, with a measured signal at $4311.9\pm 0.7$ MeV, can be observed just below the $\Sigma_{c}^{+}\bar{D}^{*0}$ threshold. From the analysis performed by the LHCb in Aaij et al. (2019), a triangle loop shown in Fig. 1, with internal lines $123$ corresponding to $D_{s}^{**-}\Lambda_{c}^{+}\bar{D}^{*0}$, can be constructed where the mass of some hypothetical $D_{s}^{**-}$ meson is set to 3288 MeV. This configuration of intermediate particles in the internal lines, with their corresponding masses, gives rise to a line shape that fit the experimental data with a peak at 4312 MeV.
In order to illustrate the ambiguity arising between a triangle mechanism and a typical resonance enhancement, we plot them simultaneously in Fig. 2. Using the Particle Data Group (PDG) values Workman et al. (2022) of the hadron masses involved in the triangle diagram, one can compute the integral in equation (2) and plot the $|I|^{2}$ contribution as a function of the invariant mass of the final state, which in this case is $J/\psi p$. A peak around the 4312 MeV can be produced by setting the values of $\epsilon=1$ MeV and $\Lambda=3000$ MeV to regulate the integral. Care must be taken in considering these parameters since this would suggest that all the particles in the internal lines are stable point hadrons which is not the case. The same argument was used by LHCb to rule out the TS interpretation for the $P^{N}_{\psi}(4312)^{+}$ state Aaij et al. (2019). The resonance interpretation can be implemented by using the Breit-Wigner parametrization to fit the experimental cross-section. The Breit-Wigner amplitude is given by
@@
where $E$ is the scattering energy, $M$ and $\Gamma$ are the mass and decay width of the unstable state. For both the Breit-Wigner and TS amplitudes, we multiply a projected phase space of the $J/\psi p$ invariant mass and some arbitrary high-order degree polynomial, following the expression
@@
where $\rho(\sqrt{s})$ is the phase space factor, $F(\sqrt{s})$ is the constructed amplitude (either Breit-Wigner or the triangle integral $I$), and $B(\sqrt{s})$ is the additional polynomial background Fernández-Ramírez et al. (2019). The $dN/d\sqrt{s}$, corresponding to the resulting intensity, achieves a better fit of the experimental data by optimizing the parameters in $F(\sqrt{s})$ and $B(\sqrt{s})$. The resulting distribution is shown in Fig. 2 where a narrow peak at around 4312 MeV is produced by both line shapes. Since both mechanisms fit within the uncertainties (error bars) of the experimental data, further analysis is required to unambiguously discriminate between the two largely different mechanisms and interpretations, despite their similar line shapes.
@@
Although this triangle mechanism fits the experimental data well, it was argued to be an unlikely origin of the enhancement due to some physical constraints Aaij et al. (2019). First, the TS description lacks an appropriate hadron-rescattering threshold, as it requires the exchange of the hypothetical $D_{s}^{**-}(3288)$ hadron in the loop. This excited meson has not yet been observed and is not listed in the PDG’s compilation of $D_{s}$ states. Second, the LHCb analysis indicated that achieving a good fit with the TS mechanism necessitates the use of unphysical parameters. Specifically, when the decay width $\epsilon$ is set to some higher and more plausible value, say 50 MeV, the enhancement from the TS is no longer prominent and falls into the background. This physical argument is the main reason why a dynamic pole-based interpretation is favored for the $P_{\psi}^{N}(4312)^{+}$ state.
@@
Figure 3: Different mechanisms for the $\Lambda_{b}^{0}\to J/\psi\,pK^{-}$ reaction that can produce a peak at 4312 MeV: a) triangle loop $\Lambda_{c}^{+}D_{s}^{**-}\bar{D}^{*0}$, b) triangle loop $J/\psi\Lambda^{*}p$, c) $P_{\psi}^{N}(4312)^{+}$ resonance, and d) coupled-channel decay.
Furthermore, there are other triangle diagrams that can produce peaks in the $J/\psi p$ invariant mass Liu et al. (2016b); Bayar et al. (2016). For instance, a $D_{s}^{**-}\Lambda_{c}^{+}\bar{D}^{0}$ triangle loop can generate an enhancement close to the $P_{\psi}^{N}(4312)^{+}$ state as well. In Bayar et al. (2016), they considered multiple diagrams where a general $\Lambda^{*}$ hyperon is assigned as particle 1 (which can couple to $K^{-}p$), a $c\bar{c}$ meson for particle 2, and $p$ for particle 3. For each $c\bar{c}$ candidate, they used equations (5) and (6) to obtain the relevant range of the hyperon mass and the range of the resulting peak. Further analysis of the branching ratios led them to argue the plausibility of each diagram. Among these, if we select the $\Lambda^{*}J/\psi p$ loop with the mass of $\Lambda^{*}$ within [2151, 2523] MeV, a peak can be achieved within the energy region of [4035, 4366] MeV. This is relevant to the $P_{\psi}^{N}(4312)^{+}$ since the signal lies within the allowed energy region of the mechanism.
@@
Upon further inspection, the last two intermediate particles in this diagram, $J/\psi p$, are similar to the two final states in the mechanism. This sort of process must follow the Schmid theorem, which states that if particles 2 and 3 in the internal triangle diagram rescatters to similar final states 2+3, then the TS will simply be absorbed by the $s$-wave of a tree-level diagram and will not produce observable contributions Schmid (1967). In other words, this mechanism may also be ruled out if any manifestation of the triangle diagram is argued to not be detectable. However, it has been shown in Anisovich and Anisovich (1995) that if one considers an inelastic scattering process in the final vertex, the Schmid theorem becomes irrelevant as the triangle may still be present under coupled-channel effects. Moreover, if a physically plausible non-zero decay width $\epsilon$ is used in the triangle, the strength of its signal is found to be comparable to that of the tree-level diagram and will also no longer be negligible Debastiani et al. (2019).
In any case, we consider both aforementioned triangle diagrams in our study as we simply aim to simulate prominent line shapes from triangle mechanisms and investigate their contributions regardless of any physical constraints. These line shapes will be used to train the DNN. One may argue that doing this allows us to possibly “trick” our machine and really test the limits of its abilities. It is up to the DNN to determine a more descriptive mechanism based solely on a pure line-shape analysis. Our benchmark is to assess if the DNN will consistently rule out these hypothetical triangle diagrams in favor of the dynamic pole structure for the $P_{\psi}^{N}(4312)^{+}$ state. Therefore, it follows that we take into account both these two triangle mechanisms in our study. From this point on, the first triangle diagram discussed will be referred to as triangle loop A, and the second as triangle loop B, as illustrated in Fig. 3(a) and 3(b), respectively.